OpenStudy (anonymous):

Will someone please help me with GCF? @satellite73

OpenStudy (anonymous):

of what?

OpenStudy (anonymous):

well, this is the first question idk how to do- "Find the GCF of the following literal terms: m^7 n^4 p^3 and mn^12 o^5"

OpenStudy (anonymous):

i meant p^5 for the last one

OpenStudy (anonymous):

i take it you mean \(m^7 n^4 p^3\) and \(mn^{12}p^5\)

OpenStudy (anonymous):

Yes, I just don't know how to type it so that it shows up like that. But yes that is what I meant.

OpenStudy (anonymous):

take each factor you see to the LOWEST power you see it

OpenStudy (anonymous):

Does that mean if I see m to the 7th power I'm supposed to change it to m to the first power?

myininaya (myininaya):

\[gcf(m^7 \cdot n^4 \cdot p^3 , m^1 \cdot n^{12} \cdot p^5 )=m^{\min(7,1)} \cdot n^{\min(4,12)} \cdot p^{\min(3,5)}\] where the min(a,b) means find the smaller of a and b and that is your power for that variable.

OpenStudy (anonymous):

no, the lowest one, not the highest you have \(m^7\) in one term, but only \(m^1\) in the second, only use \(m^1\) or just \(m\)

OpenStudy (jdoe0001):

\(\bf (m^7 n^4 p^3) ( mn^{12} o^5)\implies ({\color{red}{ m^1}}\cdot m^6 {\color{red}{ n^4}} p^3) ( {\color{red}{ m^1n^4}}\cdot n^{8} o^5)\)

OpenStudy (anonymous):

thank you I understand! but how do you do it when the powers aren't listed- "xxyyyzz and xxxxzzz"

OpenStudy (jdoe0001):

you split them up

OpenStudy (anonymous):

What does that mean

myininaya (myininaya):

You can also look at it like this: \[\gcd(x^2 y^3 z^2, x^4 y^0 z^3)=x^{\min(2,4)} y^{\min(3,0)}z^{\min(2,3)}\]

OpenStudy (jdoe0001):

\(\bf a^{\color{red}{ n}}\cdot a^{\color{red}{ m}}\cdot a^{\color{red}{ z}}\implies a^{{\color{red}{ n+m+z}}} \\ \quad \\ a^{25}\implies a^{10+10+5}\implies a^{10}a^{10}a^5\)

myininaya (myininaya):

do you know what the min(2,4)=? (in other words, which is smaller 2 or 4?)

OpenStudy (anonymous):

^ 2 is smaller

myininaya (myininaya):

\[\gcd(x^2 y^3 z^2, x^4 y^0 z^3)=\] \[x^{\min(2,4)} y^{\min(3,0)}z^{\min(2,3)}=\gcd(x^2 y^3 z^2, x^4 y^0 z^3)=x^2 y^{\min(3,0)}z^{\min(2,3)}\] So see I replace the min(2,4) with 2

myininaya (myininaya):

you try the rest

OpenStudy (anonymous):

Thank you so much everyone who is helping!! And okay since the lowest for y is 0 there is no y in the answer and since the lowest for both x and z is 2 the answer should be \[x ^{2}y ^{2}\]

myininaya (myininaya):

you meant that y to be a z right?

OpenStudy (anonymous):

oh yeah oops. I meant x squared z squared because there is no y since the lowest was 0

myininaya (myininaya):

greatness

OpenStudy (anonymous):

Yay!! Thank you! I have one last question. Can you please tell me if I worked this problem correctly? \[8x ^{6}y ^{5} - 3x ^{8}y ^{3}\] so I was gonna say the lowest between the numbers is 3, the lowest factor of x is 6, and the lowest factor of y is 2. but I'm confused because it says to subtract.

OpenStudy (anonymous):

i meant to lowest of y is 3

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

how do I write it out as an answer though

OpenStudy (jdoe0001):

\(\bf 8x ^{6}y ^{5} - 3x ^{8}y ^{3}\implies 8{\color{red}{ x^{6}}}{\color{red}{ y^3}}y ^{2} - 3x^2\cdot {\color{red}{ x^{6}y ^{3}}}\)

OpenStudy (jdoe0001):

take the GCF out and leave everything else there

OpenStudy (anonymous):

Okay thank you so much.

OpenStudy (jdoe0001):

yw